• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 23:58:31 on July 24, 2006

    Entry number 175161

    keyword=Beam Time Accounting
    Hall A

    07/24/06 1600-2400
    Experiment #: CSR
    Shift Leader: Steffen Strauch
    Shift Workers: Meryl Reyhan, Adam Sarty

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0.000.000.000.000.000.000.000.000.00
    Beam Available
    but Not in Use
    0.000.000.000.000.000.000.000.000.00
    Beam Not Available
    or Unacceptable
    1.001.001.001.001.001.001.001.008.00
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    1.000.001.001.001.001.000.001.006.00
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.001.000.000.000.000.001.000.002.00
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.001.000.000.000.000.000.000.001.00
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.000.000.000.000.000.001.000.001.00
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: 17:00-18:00 Controlled access to reset ROC3 (while MCC tuning continued). Controlled access around 22:00 to reset HRS-L Q2 quench fault, and hook up T9 triggers in HRS-L. End shift waiting for beam to return.

    Data Entry by: AJ Sarty